OpenStudy (anonymous):

A 5 foot-long cylindrical pipe has an inner diameter of 6 feet and an outer diameter of 8 feet. If the total surface area is kn(n=Pi) ,what is the value of k ?

OpenStudy (anonymous):

Area = 2.pi.r.h + 2.pi.R. h + 2[ pi(R^2 - r^2)] = 2.pi.5 (3 + 4) + 2[pi (4^2 - 3^2)] = pi.70 + 2.pi.7 = 84.pi so k = 84

OpenStudy (anonymous):

ty